On 29 August, the ship was moved in DP mode to Site 1036 and while tripping pipe a PUPPI was dropped on the locations of planned Holes 1036B and 1036C. The ship then returned to Hole 856H.


A reentry/logging bit was run because the problem was assumed to be ledges based on Leg 139 logs and open end pipe behavior. The hole was reentered in 45 min. The reentry/cleanout bit could not be worked past an obstruction at 61 m, where a ledge was evident on Leg 139 logs. The logging bit was pulled and severe wear was noted on the throat and outside of the bit face, indicating a hard material had blocked the hole. The hole was unstable and bridging over so a 9-7/8" Smith three-cone drill bit was run to clean out the hole. The hole was reamed as required from 36.5 to 93.5, 126.5 to 209.5, 375.5 to 395.5, and 487.5 to 500.0 mbsf. A conditioning trip was made back to 85 m, and 11 m of fill was cleaned out on bottom. Sepiolite mud (115 bbl) was spotted in the hole in an effort to stabilize it and reduce the inflow of hot water and the bit was dropped in the bottom of the hole. The open end pipe was pulled to 85 m for logging. The Triple Combo log was run to 495 m. The GR/Sonic/FMS log was run to 491 m; however, the FMS data was erratic (the internal tool temperature read 86°C). The tool appeared to be slipping with and without the wireline heave compensator. The tool was pulled, and an external temperature tab read 182°C (the FMS temperature limit is 175°C). The Schlumberger FMS logging tool has an internal temperature sensor that was reading 90°C. The FMS sonde was damaged by heat and the transmitter was damaged, which left only one operational FMS tool. The open end drill string was run to 500 m, and the hole was circulated 1 hour (2 hole volumes) to cool the hole. The pipe was pulled back to 85 m, and the second run was made with the GR/FMS (no Sonic) to 385 m. The second FMS run was successful, but the external temperature tab read 176°C when the tool was retrieved, so the geochemical log was canceled because its temperature limit is 130°C. The pipe was pulled, clearing the rotary table at 1330 hr on 1 September. The primary beacon was generating erratic signals, so it was recovered. The backup beacon was turned off for the return to Site 1035. The pipe cleared the rotary at 1330 hr on 1 September.

To 169 Return To Hole 858G

To 169 Table of Contents

Publications Homepage

ODP Homepage